OpenStudy (orion1):

A whole seller supplies 90 books while he was ordered to send 80 books by a retailer. If 15%trade discount was given, what was the net discount to the retailer? Please help me with the solution and answer is required in percent. The book says answer is 27.5% but I am looking way from getting that. Please help me

OpenStudy (orion1):

Please helpASAP

OpenStudy (mww):

Here you need to add the discount from extra books to the original 15% since he got 10 extra books more than the 80 quota, he is already given 10/80 = 1/8 = 12.5% discount. add to the 15% and you get your answer.